# Calibration Attention: Instance-wise Temperature Scaling for Vision Transformers

### v1

@@ Abstract
Probability calibration is critical when Vision Transformers (ViTs) are deployed in risk-sensitive applications. However, the conventional solution—post-hoc temperature scaling—employs a single global scalar and requires tuning on a separate validation set. We propose Calibration Attention (CalAttn), a novel drop-in module designed to dynamically learn an adaptive, per-instance temperature directly from the ViT’s CLS token. Empirical evaluations on CIFAR-10/100, MNIST, TinyImageNet, ImageNet1K demonstrate that CalAttn substantially reduces calibration error—by up to a factor of 4—in models such as ViT-224, DeiT, and Swin Transformers, while introducing fewer than 0.1% additional parameters. Notably, the learned temperatures exhibit tight clustering around 1.0, starkly contrasting the large global values typically needed by standard temperature scaling methods. Calibration Attention offers a straightforward, efficient, and architecture-agnostic strategy for producing trustworthy and well-calibrated probabilities in state-of-the-art Vision Transformers.
@@ 1 Introduction
A classifier is well-calibrated when its predicted probabilities reflect the true likelihood of correctness: among all predictions made with 80% confidence, roughly 80% should in fact be correct. Calibration is crucial whenever confidence guides downstream decisions—especially in risk-sensitive domains such as medical diagnosis Mehrtash et al. [2020], autonomous driving Feng et al. [2021], and financial trading Liang et al. [2024a].
Confidence matters. Although modern deep neural networks—including both CNNs and Vision Transformers (ViTs)—achieve state-of-the-art accuracy, they are often mis-calibrated Guo et al. [2017], Minderer et al. [2021]. The common remedy is temperature scaling, a post-hoc technique that multiplies logits by a single global scalar $T^{\star}$ fitted on a validation set Guo et al. [2017]. While effective, this approach (i) assumes uniform difficulty across inputs, ignoring sample-specific uncertainty, and (ii) is decoupled from training, so the backbone never receives gradients related to calibration.
Despite their excellent accuracy on large-scale benchmarks, ViTs (ViT Dosovitskiy et al. [2021], DeiT Touvron et al. [2021], Swin Liu et al. [2021]) suffer from the same calibration issues as CNNs. Early reports hinted that ViTs might be “naturally” calibrated Touvron et al. [2021], Ridnik et al. [2023], but our systematic re-evaluation (Section 3) shows otherwise. After optimal global temperature scaling, the calibration gap between CNNs and ViTs essentially disappears, echoing Minderer et al. [2021]. This underscores the two limitations of global scaling mentioned above.
A hidden cue to learn per-sample uncertainty. While prior works have speculated that ViTs may encode uncertainty cues [Ridnik et al., 2023], we provide the first quantitative evidence that the $\ell_{2}$ norm of the final [CLS] token embedding, $\lVert\mathbf{z}_{\mathrm{CLS}}\rVert_{2}$, is moderately correlated with model confidence (Pearson $r=0.45$; Fig. 2). Larger norms correspond to easier, over-confident samples, whereas smaller norms flag harder ones. Leveraging this cue, we introduce Calibration Attention (CalAttn)—a two-layer MLP (<0.1% parameters) that maps $\mathbf{z}_{\mathrm{CLS}}$ to a positive temperature $s(\mathbf{z})$. CalAttn scales logits during both training and inference and is trained jointly with the backbone using cross-entropy plus a small Brier penalty. It therefore (i) adapts confidence to individual samples and (ii) integrates calibration into the training objective.
CalAttn in training. Prior methods largely rely on global temperature scaling [Guo et al., 2017], which overlooks per-sample uncertainty and disconnects calibration from training. ViT-Calib [Ridnik et al., 2023] extends calibration to Transformers but lacks mechanisms to exploit internal difficulty signals. CalAttn addresses both limitations by leveraging the predictive value of the [CLS] token norm. A lightweight two-layer MLP predicts an instance-specific temperature $s(\mathbf{z}_{\mathrm{CLS}})$, used to scale logits during training and inference. The head is trained end-to-end with the backbone using cross-entropy and a Brier penalty, enabling dynamic self-calibration with minimal overhead.
The contributions of this paper are summarized as follows:
  1. 1.
    We benchmark calibration performance across CNNs and ViTs using seven calibration losses on CIFAR-10/100 and Tiny-ImageNet, showing that ViTs do not inherently offer better calibration after tuning.
  2. 2.
    We propose $\|z_{\text{CLS}}\|_{2}$ as an intrinsic difficulty indicator and use it to drive instance-wise temperature scaling.
  3. 3.
    We introduce CalAttn, the first ViT-compatible, train-time calibration module that achieves state-of-the-art reliability with negligible overhead.
@@ 2 Method
To address the above limitations, our aim is to replace the single, global temperature used in post-hoc scaling with an instance-wise, image-adaptive temperature that is learned jointly with model parameters. Fig. 1 presents an overview of our approach, and full notation is provided in Appendix A.
Figure 1: ViT +CalibrationAttention (CalAttn). A standard Vision Transformer pipeline is augmented with a lightweight Calibration Head (FC + GELU + FC + Softplus) that reads the final CLS embedding and predicts a strictly–positive, per-sample temperature $s(\mathbf{z})$. The class logits produced by the usual classifier head are divided by this scale before the Soft-max, enabling image-adaptive cooling or sharpening of the output distribution. The main backbone mirrors the architecture of Dosovitskiy et al. [2021]; the tables on the right are supposed a 3-class prediction task.
@@ 2.1 Preliminaries
Given an RGB image $\mathbf{x}\!\in\!\mathbb{R}^{3\times H\times W}$, a ViT prepends a learned [CLS] token and produces, after $L$ encoder blocks, a global embedding $\mathbf{z}_{\text{CLS}}\!\in\!\mathbb{R}^{d}$. A linear head  $\boldsymbol{\ell}=\mathbf{W}_{\text{cls}}\mathbf{z}_{\text{CLS}}+b_{\text{cls}}$ yields class-logits $\boldsymbol{\ell}\in\mathbb{R}^{C}$, which are turned into probabilities by a soft-max with temperature $T$,
$\sigma_{i}(\boldsymbol{\ell},T)\;=\;\frac{\exp(\ell_{i}/T)}{\sum_{j=1}^{C}\exp(\ell_{j}/T)}.$ (1)
Limitations of global scaling.
Temperature scaling [Guo et al., 2017] fits a single scalar $T^{\star}$ on a held-out set and deploys $\hat{y}_{i}=\sigma_{i}(\boldsymbol{\ell},T^{\star})$ at test time. Because $T^{\star}$ is constant, it cannot reconcile the heteroscedastic [^1] Easy images lead to large $\|\mathbf{z}_{\text{CLS}}\|_{2}$ and over-confidence; difficult images do the opposite. confidence of individual samples, and the backbone never sees calibration gradients.
2.2 Calibration Attention (CalAttn)
A per-sample temperature head.
After the Transformer encoder, the first token yields a global classification embedding $\mathbf{z}\in\mathbb{R}^{d}$, commonly known as the CLS token. Calibration Attention (CalAttn) maps this embedding through a lightweight two-layer MLP to predict an adaptive, per-sample temperature:
$\boxed{s(\mathbf{z})=\operatorname{softplus}\left(\mathbf{w}_{2}^{\top}\operatorname{GELU}(\mathbf{W}_{1}\mathbf{z})+b_{2}\right)+\varepsilon}$ (2)
where $\mathbf{W}_{1}\in\mathbb{R}^{h\times d}$, $\mathbf{w}_{2}\in\mathbb{R}^{h}$, $b_{2}\in\mathbb{R}$, and $\varepsilon=10^{-6}$. Parameters are initialized as $\mathbf{w}_{2}\leftarrow\mathbf{0}$ and $b_{2}\leftarrow\ln(e^{1}-1)$, ensuring $\operatorname{softplus}(b_{2})\approx 1$. Thus, CalAttn initially matches the baseline model and deviates only when driven by calibration improvements.
The final calibrated probability distribution is computed as:
$\widehat{\boldsymbol{y}}=\operatorname{softmax}\left(\frac{\boldsymbol{\ell}}{s(\mathbf{z}_{\text{CLS}})}\right),$ (3)
where scaling by the adaptive scalar $s(\mathbf{z})$ preserves the original class ranking, dynamically cooling overly confident predictions ($s>1$) or sharpening under-confident ones ($s<1$).
MLP Design.
The choice of a small two-layer MLP for CalAttn is guided by several key insights:
  1. 1.
    Implicit Difficulty Signal in $\mathbf{z}$. Empirically, the magnitude and geometry of the CLS embedding encode information about sample difficulty. CalAttn directly leverages this intrinsic representation to predict adaptive calibration scales.
  2. 2.
    Expressivity with Minimal Cost. A single hidden-layer MLP with GELU activation universally approximates scalar functions. With a modest width $h=128$, this module adds only $hd+d+h+1$ parameters, equivalent to less than $0.1\%$ of the parameters in typical Transformers (e.g., DeiT-Small), ensuring negligible computational overhead.
  3. 3.
    Smooth and Positive Scaling via Softplus. Softplus ensures strictly positive scaling ($s(\mathbf{z})>0$), provides smooth and stable gradients, and avoids numerical instabilities common with non-smooth alternatives like ReLU [Glorot et al., 2011]. The small offset $\varepsilon$ further guarantees it.
  4. 4.
    Calibration-Aligned Gradient Dynamics. For a combined cross-entropy and Brier loss (Eq. 5), the gradient w.r.t. the scaling parameter is 222where $\hat{c}=\arg\max_{i}\ell_{i}$ is the predicted class, $\hat{y}_{i}=\exp(\ell_{i}/s)/\sum_{j}\exp(\ell_{j}/s)$ the soft-max probability for class $i$, $\mathbf{1}[y=\hat{c}]=1$ if the true label $y$ equals $\hat{c}$ and 0 otherwise, and $\ell_{j}$ denotes the $j$-th logit before scaling.:
    $\frac{\partial\mathcal{L}}{\partial s}=(\hat{y}_{\hat{c}}-\mathbf{1}[y=\hat{c}])\Bigl{(}\ell_{\hat{c}}-\sum_{j}\hat{y}_{j}\ell_{j}\Bigr{)}\,s^{-1}.$ (4)
    This gradient term is positive when predictions are incorrect yet overly confident (thus increasing $s$), and negative when correct but underconfident (decreasing $s$), directly aligning confidence with accuracy to reduce calibration error.
Heteroscedastic-Noise Perspective.
CalAttn can also be justified statistically. Assume raw logits scale inversely with a latent, per-image noise level $\sigma(x)$, i.e., $\boldsymbol{\ell}\propto 1/\sigma(x)$. The Bayes-optimal softmax would thus use temperature scaling $T(x)\propto\sigma(x)$.
CalAttn achieves this ideal scenario333“ideal” mean the Bayes-optimal, instance-wise temperature $T^{\star}(x)\propto\sigma(x)$, where $\sigma(x)$ is the latent per-sample noise scale: this unique scalar exactly aligns a model’s predicted confidence with its true accuracy. end-to-end by optimizing a differentiable calibration-aware objective:
$\mathcal{L}=\mathcal{L}_{\mathrm{CE}}+\lambda\,\bigl{\|}\widehat{\boldsymbol{y}}-\mathbf{e}_{y}\bigr{\|}_{2}^{2},$ (5)
where $\mathbf{e}_{y}$ is the one-hot ground-truth vector and $\lambda=0.1$. Also, the Brier term serves as a differentiable proxy for Expected Calibration Error (ECE) [Liang et al., 2024b, Blasiok et al., 2023]. As the only scalar that can modify confidence while preserving class ranks, the learned temperature $s(\mathbf{z}_{\text{CLS}})$ naturally converges toward the optimal temperature (Proof in Appendix E):
$\nabla_{s}\mathcal{L}=0\quad\Longrightarrow\quad s(\mathbf{z}_{\text{CLS}})\approx T(x)_{\text{optimal}}.$ (6)
Thus, CalAttn effectively replaces traditional two-stage post-hoc temperature searches with a one-stage, fully differentiable calibration solution as depicted in Table 4.
2.3 CLS token is a good “Thermometer”
Global summary token. In ViT–style architectures a class token is prepended to the patch sequence and attends to all patches at every layer. After $L$ encoder blocks the first-token embedding $\mathbf{z}_{\text{CLS}}\in\mathbb{R}^{d}$ is a compact, information-rich summary: a linear head produces the logits $\boldsymbol{\ell}=\mathbf{W}\mathbf{z}_{\text{CLS}}.$ Empirically, the norm and direction of $\mathbf{z}_{\text{CLS}}$ correlate with (i) image difficulty, (ii) inter-class margin and (iii) distribution shift [Naseer et al., 2021, Zhou et al., 2022, Minderer et al., 2021].
Logit scale and confidence. For a linear classifier the magnitude of the pooled token is a proxy for the scale of the logits:
$\bigl{\|}\mathbf{z}_{\textsc{cls}}\bigr{\|}_{2}\;\;\xrightarrow{\;\mathbf{W}\,\cdot\;}\;\;\text{larger logits}\;\;\xrightarrow{\;\sigma(\,\cdot\,)\;}\;\;\text{higher soft-max confidence}.$
Empirically (Figure 2 plots the L2 norm $\|\mathbf{z}_{\text{CLS}}\|_{2}$ against soft-max confidence for CIFAR-10/100.) we observe a clear, monotonic trend: Easy images with large $\|\mathbf{z}_{\textsc{cls}}\|_{2}$ tend to be over-confident; while hard images with small $\|\mathbf{z}_{\textsc{cls}}\|_{2}$ tend to be under-confident. A single post-hoc temperature $T^{\star}$ inevitably calibrates only a narrow slice of this spectrum, leaving both tails mis-calibrated. The moderate monotonic trend (Pearson r=0.45r{=}0.45, Spearman $\rho{=}0.45$) confirms that ViTs already encode sample difficulty; CalAttn exploits this signal through Eq. 2.
Figure 2: CLS-norm versus soft-max confidence on CIFAR-10/100 with ViT-224. Blue: individual images; red: mean in 15 equal-width bins. The positive correlation shows that $\|\mathbf{z}_{\text{CLS}}\|_{2}$ diagnoses sample difficulty, which CalAttn converts into an adaptive temperature.
2.4 Temperature Inference
At test time we run one forward pass: compute $s(\mathbf{z})$, divide logits, apply soft-max. No extra tuning or ensemble passes are required (See Alogorithm 1). CalAttn’s simplicity, tiny footprint and substantial drops in ECE (Table 3) make it a practical drop-in module for trustworthy Vision Transformers.
Algorithm 1 Per-sample Temperature Scaling with Calibration Attention
1:CLS token $\mathbf{z}_{\mathrm{cls}}\!\in\!\mathbb{R}^{d}$, logits $\boldsymbol{\ell}\!\in\!\mathbb{R}^{C}$, weights $W_{1}\!\in\!\mathbb{R}^{h\times d}$, $W_{2}\!\in\!\mathbb{R}^{1\times h}$, bias $b_{2}\!=\!\ln(e^{1}-1)$ $\triangleright$ $\text{Softplus}(b_{2})\!\approx\!1$
2:Calibrated class-probabilities $\hat{\mathbf{y}}\!\in\!\mathbb{R}^{C}$

### v2

# Calibration Attention: Learning Reliability-Aware Representations for Vision Transformers

@@ Abstract
Most calibration methods operate at the logit level, implicitly assuming that miscalibration can be corrected without changing the underlying representation. We challenge this assumption and propose Calibration Attention (CalAttn), a representation-aware calibration module for vision transformers that couples instance-wise temperature scaling to transformer token geometry under a proper scoring objective. CalAttn predicts a sample-specific temperature from the [CLS] token and backpropagates calibration gradients into the backbone, thereby reshaping the uncertainty structure of the representation rather than post-hoc adjusting confidence. This yields token-conditioned uncertainty modulation with negligible overhead ($0.1\%$ additional parameters). Across multiple datasets with ViT/DeiT/Swin backbones, CalAttn consistently improves calibration while preserving accuracy, achieving relative ECE reductions of $3.7\%$ to $77.7\%$ over strong baselines across diverse training objectives. Our results indicate that treating calibration as a representation-level problem is a practical and effective direction for trustworthy uncertainty estimation in transformers.
Figure 1: Vision Transformer with Calibration Attention. A lightweight calibration head reads the final [CLS] token and predicts a strictly positive, per-sample temperature $s(\mathbf{z})$. The class logits produced by the standard classifier head are divided by this scale prior to the softmax, enabling representation-conditioned cooling or sharpening of predictive confidence. CalAttn learns reliability-aware representations by conditioning confidence on the [CLS] embedding, replacing global post-hoc temperature scaling.
@@ 1 Introduction
A probabilistic classifier is well calibrated when its predicted confidence matches empirical correctness: among predictions assigned confidence $p$, roughly a $p$ fraction are correct. Calibration is therefore central to decision-making under risk and is consequential in safety-critical settings such as medical diagnosis, autonomous driving, and finance Mehrtash et al. (2020); Feng et al. (2021); Liang et al. (2024b).
Limitations of logit-level calibration.
Despite strong accuracy, modern deep networks—including Vision Transformers (ViTs)—are often miscalibrated Guo et al. (2017); Minderer et al. (2021); Ovadia et al. (2019). The canonical remedy, temperature scaling (TS), fits a single scalar temperature on a validation set and rescales logits at test time. While simple and effective in-distribution, TS implicitly assumes that miscalibration is (i) homogeneous across inputs and (ii) expressible as a logit-only correction. In practice, predictive uncertainty is heterogeneous and distribution-dependent: different samples exhibit different ambiguity, and shifts in data distribution can substantially change uncertainty patterns Ovadia et al. (2019). Consequently, a global temperature can over-dampen genuinely confident predictions while under-correcting ambiguous ones, producing brittle calibration under distribution shift even when accuracy remains high. Training-time approaches (e.g., label smoothing, focal-style objectives, and differentiable calibration surrogates) mitigate miscalibration but typically rely on static hyperparameters and do not explicitly condition uncertainty on the learned representation Mukhoti et al. (2020a); Müller et al. (2019); Tao et al. (2023); Liang et al. (2024a).
Representation geometry as a source of uncertainty.
We argue that, in transformers, miscalibration is not solely an artifact of the output layer but is tied to the geometry of the learned representation. ViTs aggregate global evidence into the [CLS] token, which summarizes patch information and encodes signals related to class separability, sample ambiguity, and distributional irregularities. These signals are largely invisible to post-hoc logit transformations. This motivates the view that calibration should be conditioned on internal token geometry, rather than treated purely as a logit correction.
Representation-aware calibration.
Motivated by this perspective, we introduce representation-aware calibration: an end-to-end paradigm in which uncertainty modulation is modeled as a function of internal token representations. Concretely, we propose Calibration Attention (CalAttn), a lightweight two-layer MLP that predicts an instance-specific temperature from the [CLS] embedding and rescales logits accordingly:
$T(x)=g_{\phi}(\mathrm{CLS}_{\theta}(x)),$
where $g_{\phi}$ is trained jointly with backbone parameters $\theta$. This design couples calibration with representation learning: calibration gradients propagate through the [CLS] channel, enabling the backbone to adjust its uncertainty geometry during training. In contrast to adaptive temperature scaling variants that regress temperatures from logits or apply post-hoc corrections on frozen features, CalAttn explicitly links token learning dynamics with uncertainty estimation in ViTs.
Contributions.
Our contributions are:
  • Paradigm. We formalize representation-aware calibration for transformers, where uncertainty modulation is conditioned on internal token geometry rather than applied as a purely logit-level post-hoc transform.
  • Method. We propose Calibration Attention (CalAttn), a minimal module that predicts sample-specific temperatures from [CLS] and injects calibration gradients into representation via a proper scoring objective.
  • Evidence. Across MNIST, CIFAR-10/100, Tiny-ImageNet, and ImageNet-1K, and across ViT/DeiT/Swin backbones, CalAttn consistently improves calibration while preserving accuracy and remains complementary to standard post-hoc methods.
@@ 2 Method
We aim to replace post-hoc global temperature scaling with an instance-wise, representation-conditioned temperature that is learned jointly with the backbone. Given an input image, a ViT produces a final-layer [CLS] embedding that summarizes global evidence. CalAttn attaches a lightweight temperature head to this embedding, predicts a strictly positive scale for each sample, rescales logits before the softmax, and trains end-to-end with a calibration-aware proper scoring objective. This coupling ensures that calibration gradients propagate through the [CLS] pathway and can reshape the representation during training, rather than only correcting confidence after the fact. Figure 1 provides an overview; notation is summarized in Appendix A.
@@ 2.1 Preliminaries
Given an input image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$, a Vision Transformer prepends a learned [CLS] token and, after $L$ encoder blocks, outputs the final [CLS] embedding $\mathbf{z}_{\text{CLS}}\in\mathbb{R}^{d}$. A linear classifier head produces logits
$\boldsymbol{\ell}=\mathbf{W}_{\text{cls}}\mathbf{z}_{\text{CLS}}+b_{\text{cls}}\in\mathbb{R}^{C},$
which define probabilities via the temperature-scaled softmax
$\sigma_{i}(\boldsymbol{\ell},T)=\frac{\exp(\ell_{i}/T)}{\sum_{j=1}^{C}\exp(\ell_{j}/T)}.$
Global temperature scaling.
Temperature scaling (TS) fits a single scalar $T^{\star}$ on a validation set and applies it at test time Guo et al. (2017). This procedure is decoupled from representation learning and cannot condition confidence adjustment on sample-specific uncertainty. Under heterogeneous uncertainty or distribution shift, a single global temperature may under-correct ambiguous samples while over-smoothing confident ones.
2.2 Calibration Attention
Instance-wise, representation-conditioned temperature.
Let $\mathbf{z}\triangleq\mathbf{z}_{\text{CLS}}$ denote the final [CLS] embedding. CalAttn predicts each scale (temperature) via a two-layer MLP:
$s(\mathbf{z})=\mathrm{Softplus}\!\left(\mathbf{w}_{2}^{\top}\mathrm{GELU}(\mathbf{W}_{1}\mathbf{z})+b_{2}\right)+\varepsilon,$ (1)
where $\mathbf{W}_{1}\in\mathbb{R}^{h\times d}$, $\mathbf{w}_{2}\in\mathbb{R}^{h}$, $b_{2}\in\mathbb{R}$, and $\varepsilon=10^{-6}$ ensures strict positivity. We use an identity-preserving initialization with $\mathbf{w}_{2}=0$ and $b_{2}=\ln(e^{1}-1)$ such that $s(\mathbf{z})\approx 1$ at the start of training, making CalAttn initially equivalent to the baseline model. The calibrated predictive distribution is
$\widehat{\boldsymbol{y}}=\mathrm{softmax}\!\left(\boldsymbol{\ell}/s(\mathbf{z})\right).$ (2)
Since scaling preserves logit ordering, CalAttn primarily modulates confidence rather than class ranking.
Calibration-aware proper scoring objective.
We train the backbone and CalAttn jointly by minimizing a proper scoring objective consisting of cross-entropy and a Brier term:
$\mathcal{L}=\mathcal{L}_{\mathrm{CE}}(\widehat{\boldsymbol{y}},y)+\lambda\,\|\widehat{\boldsymbol{y}}-\mathbf{e}_{y}\|_{2}^{2},$ (3)
where $\mathbf{e}_{y}$ is the one-hot label and $\lambda=0.1$ unless stated otherwise. The Brier term acts as a differentiable surrogate encouraging calibrated probabilities, and its gradients backpropagate through $s(\mathbf{z})$ into the [CLS] representation (see Appendix E).
Gradient signal through the temperature head.
Differentiating $\mathcal{L}$ with respect to the (positive) scale $s$ yields
$\frac{\partial\mathcal{L}}{\partial s}$ $=\underbrace{\frac{\ell_{y}-\sum_{j}\hat{y}_{j}\ell_{j}}{s^{2}}}_{\text{CE term}}\;+\;\underbrace{\lambda\frac{\partial}{\partial s}\|\widehat{\boldsymbol{y}}-\mathbf{e}_{y}\|_{2}^{2}}_{\text{Brier term}}.$ (4)
Using the softmax Jacobian $J(\widehat{\boldsymbol{y}})=\mathrm{diag}(\widehat{\boldsymbol{y}})-\widehat{\boldsymbol{y}}\widehat{\boldsymbol{y}}^{\top}$, the Brier contribution can be written as
[…]
2.3 [CLS] Token as an Uncertainty Carrier
The [CLS] token aggregates global context across patches and layers. Prior work suggests that its geometry encodes signals related to separability, sample difficulty, and robustness Naseer et al. (2021); Zhou et al. (2022); Minderer et al. (2021). CalAttn leverages the full embedding to learn a non-linear mapping from representation to uncertainty modulation, rather than relying on a single handcrafted statistic. At inference, CalAttn adds a single lightweight head: compute $s(\mathbf{z})$, rescale logits, and apply softmax (Algorithm 1).
Figure 2: [CLS] $\ell_{2}$-norm versus maximum softmax probability on CIFAR-10/100. Blue dots are samples; red curves show bin means. Correlation is moderate and non-deterministic, motivating learning a representation-aware calibration function.
Algorithm 1 Per-sample Temperature Scaling with CalAttn
Input: [CLS] token $\mathbf{z}_{\mathrm{cls}}\!\in\!\mathbb{R}^{d}$, logits $\boldsymbol{\ell}\!\in\!\mathbb{R}^{C}$
Parameters: $W_{1}\!\in\!\mathbb{R}^{h\times d}$, $W_{2}\!\in\!\mathbb{R}^{1\times h}$, $b_{2}=\ln(e^{1}-1)$, $\varepsilon=10^{-6}$
Output:
Calibrated probabilities $\hat{\mathbf{y}}\!\in\!\mathbb{R}^{C}$